# DPMambaIR: All-in-One Image Restoration via Degradation-Aware Prompt State Space Model

### v2

@@ V Conclusion
We propose DPMambaIR, a novel All-in-One framework for image restoration capable of handling diverse degradation types. The core of our framework is a Degradation-Aware Prompt State Space Model (DP-SSM) that leverages a fine-grained degradation extractor. This design enables the dynamic integration of degradation features into the state-space modeling process, allowing for adaptive handling of complex degradation scenarios while maintaining global degradation awareness. Furthermore, a lightweight High-frequency Enhancement Block (HEB) was introduced to complement the main framework by enhancing high-frequency details restoration with negligible computational overhead. This study underscores the importance of fine-grained degradation modeling and dynamic feature modulation in advancing All-in-One image restoration frameworks. Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, establishing it as a promising approach for unified image restoration.

### v3

@@ V Conclusion
We propose DPMambaIR, a novel All-in-One framework for image restoration capable of handling diverse degradation types. The core of our framework is a Degradation-Aware Prompt State Space Model (DP-SSM) that leverages a fine-grained degradation extractor. This design enables the dynamic integration of degradation features into the state-space modeling process, allowing for adaptive handling of complex degradation scenarios while maintaining global degradation awareness. Furthermore, a lightweight High-frequency Enhancement Block (HEB) is introduced to complement the main framework by enhancing high-frequency detail restoration with negligible computational overhead. This study underscores the importance of fine-grained degradation modeling and dynamic feature modulation in advancing All-in-One image restoration frameworks. Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, establishing it as a promising approach for unified image restoration.